OpenStudy (anonymous):

Which property is demonstrated by this equation? (1 + 2) + 3 = 1 + (2 + 3) Commutative property Associative property Distributive property

OpenStudy (anonymous):

Which property is demonstrated by this equation? 2 + 3 = 3 + 2 Commutative property Associative property Distributive property

OpenStudy (anonymous):

Which property is demonstrated by this equation? 3(1 + 2) = (1 + 2)3 Commutative property Associative property Distributive property

OpenStudy (loser66):

1/ associative 2/commutative 3/commutative Thanks for asking, hehehe. I have to learn the name of the property too, but to me, it's tedious